OpenStudy (anonymous):

If x - 3 = 12, then what is the value of x - 3 + 3? Why?

OpenStudy (anonymous):

x - 3 + 3 = ? x - 3 = 12 I wrote it like that because it makes it easy to see how we can just replace x - 3 with 12. The second equation says that x - 3 is the EXACT SAME THING as 12, so wherever we see x - 3, we can just write 12. (keep in mind this is because we're allowed to put parentheses in this situation) x - 3 + 3 = ? 12 + 3 = ? 15 Alternatively, you could have solved for x and plugged it back into the equation. No difference, really.